### v2

@@ 5.4 Metropolitan Areas
We estimate our model specifically for residents in metropolitan areas to address concerns about potential differences between urban and rural populations. By focusing on metro areas, where labor markets may function differently due to higher population density and greater economic diversity, we aim to test whether the effects of anti-discrimination laws vary by geographic location.
Table 8 shows that when the sample is restricted to metropolitan areas, the overall patterns for men in same‑sex partnerships remain consistent, though the estimated effects of anti‑discrimination laws on labor supply are slightly smaller than in the full sample. This may suggest that the urban labor market environment, with its higher baseline employment levels and different institutional dynamics, somewhat tempers the positive impact seen elsewhere. In contrast, for women in same‑sex partnerships, the estimates become significantly negative across several outcomes, labor force participation, employment, and weeks worked. One plausible explanation is that ENDAs may improve employment stability in urban labor markets, and with high childcare costs in metro areas, a stable job by one partner might encourage women to opt out of the workforce. With a secure income from the other partner, some women may choose to stay home to manage childcare responsibilities rather than work additional hours or remain continuously employed. This pattern highlights clear geographic heterogeneity, as the dynamics in urban settings differ markedly from those in the broader sample for women in same‑sex partnerships.
@@ 5.5 Reporting Changes
Prior research has highlighted higher rates of misclassification errors before 2008, where different-sex couples were sometimes incorrectly identified as same-sex couples due to typos in the sex variable. In response, and considering the 2008 ACS redesign aimed at reducing such errors (O’Connell and Lofquist (2009)), we restrict our sample to 2008–2019 to test the robustness of our findings reported in Table 4. Table 9 presents the estimates based on this restricted sample.
@@
For men in same-sex partnerships, the results remain qualitatively similar to our full sample analysis. The estimated impacts on both labor force participation and employment continue to be positive and statistically significant, with coefficients that are comparable in magnitude to those in Table 4. Among women, the extensive margin estimates remain small and statistically insignificant, while the intensive margin results are largely consistent with the full sample, albeit with some variations in magnitude.
@@
These findings suggest that potential misclassification errors in earlier years do not drive our main conclusions. The consistency of the results between the full sample (2005–2019) and the restricted sample (2008–2019) reinforces the robustness of our empirical strategy and the validity of our conclusions regarding the impact of anti-discrimination laws on labor market outcomes for same-sex couples.
5.6 Weighting
Although the ACS recommends applying sample weights to ensure representativeness, we conduct a robustness check by re‐estimating our models without these weights, following Solon et al. (2015). This approach allows us to verify that our main results are not driven solely by the weighting scheme.
The results in Table 10 indicate that the overall impacts of anti‑discrimination laws on labor supply remain robust even when ACS weights are removed. For men in same‑sex partnerships, the magnitude of the estimated effects on both labor force participation and employment is slightly smaller compared to the weighted estimates, but they continue to be statistically significant. This suggests that the positive impact of these laws for men is not driven solely by the weighting procedure. For women, the pattern of the results is similar to the weighted sample, with some variations in magnitude across different labor outcomes.
5.7 Placebo Tests and Alternative Estimators
We next conduct a series of placebo tests and apply alternative estimators to assess the impact of anti‑discrimination laws on labor market outcomes among both same‑sex and different‑sex couples. Specifically, we utilize four distinct methods to assess the impact of anti‑discrimination laws on labor market outcomes among same‑sex and different‑sex couples: TWFE, Callaway and Sant’Anna (2021), Sun and Abraham (2021), and Borusyak et al. (2021). As a placebo test, we apply the DiD framework to different‑sex couples, who are not the target of these laws, and thus we expect no significant changes in their labor market outcomes. Panel (a) of Tables 11 and 12 report the DiD estimates for different‑sex couples, while Panels (a) and (c) of Figures 6 and 7 display the corresponding event study plots.
Next, Panel (b) of Tables 11 and 12 presents the impact of anti‑discrimination laws on same‑sex couples, while Panels (b) and (d) of Figures 6 and 7 display the corresponding event studies. In these analyses, we compare labor market outcomes for same‑sex couples in regions that have enacted the laws with those in regions that have not. The results reveal that male same‑sex couples in areas with anti‑discrimination laws experience significant improvements in labor market outcomes, particularly through increased labor force participation and higher employment rates. For female same‑sex couples, however, we observe no significant effects on the extensive margin; instead, the positive and significant impact is confined to the intensive margin, as evidenced by an increase in average weekly hours worked. Hence, our alternative estimators yield similar point estimates and comparable confidence intervals, reinforcing the robustness of our findings.
@@ 6.1 Differences in Response by Sex
The results reveal that anti-discrimination laws affect household labor allocation differently for same-sex couples by gender. Men in male same-sex partnerships experience a reduction in their labor supply gap relative to men in different-sex partnerships, while there are no significant average effects for women in female same-sex partnerships.
One potential explanation is that changes in job security, stemming from reduced discrimination, alter the household’s allocation of labor between market and non‑market activities. As documented by Badgett (2006), reduced discrimination in the labor market improves employment prospects and job security for sexual minorities. With greater job security for the primary earner, the secondary earner may reduce market participation in favor of increased household production, including child rearing. This pattern of adjustment is consistent with models of household bargaining and collective decision‑making (see, e.g., Chiappori (1992) and Lundberg and Pollak (1993)), which predict that improvements in labor market conditions for one partner can induce a reallocation of tasks within the household. In this framework, the observed increase in intra‑household differences in hours worked may reflect a shift toward greater specialization, with one partner focusing more on market production while the other devotes more time to household responsibilities. Thus, the differential responses between male and female same‑sex partnerships could arise from distinct household preferences and constraints.
Households consisting of women in same-sex partnerships typically have more children than those consisting of men in same-sex partnerships. Given this difference, it would be unsurprising to see female same-sex households shift more toward a specialization of household labor compared to male same-sex households. This difference in specialization could help explain why anti-discrimination laws have different effects on women in same-sex partnerships and men in same-sex partnerships.
To test the different responses by men and women in same-sex partnerships to anti-discrimination laws, we implement an alternative difference-in-difference model. Instead of comparing those in same-sex partnerships to those in different-sex partnerships, we compare men and women in same-sex partnerships to examine their differential responses at the household level. Do women in same-sex partnerships change their household division of labor relative to men in same-sex partnerships following anti-discrimination laws? To answer this question, we collapse the data to the household level and examine households instead of individuals. Specifically, we estimate this model:
@@
$Y_{jct}$ $=\alpha_{0}\,ENDA_{ct}+\alpha_{1}\,FemSSP_{j}+\alpha_{2}\,\bigl{(}ENDA_{ct}\times FemSSP_{j}\bigr{)}+\mu_{c}+\delta_{t}+\epsilon_{jct}$ (3)
The notation is the same as before with two main differences. First, our interaction term uses an indicator for whether household $j$ is a female same‐sex partnership, denoted by $FemSSP_{j}$. Second, $Y_{jct}$ now represents outcomes measured at the household level. The coefficient $\alpha_{2}$ is our parameter of interest and captures how female same‐sex households differ in their response to anti‑discrimination laws relative to male same‑sex households. The first outcome examined is an indicator for whether the household is a one‑earner family. We also analyze the intra‑household division of labor by measuring the absolute difference in weekly hours worked between partners, thereby capturing both the extensive and intensive margins of labor supply. Finally, we explore the potential effects on child rearing by examining both the presence of children in the household and the number of children present.
We present the results in Table 13. Anti-discrimination laws do not affect the likelihood of same-sex partnerships becoming a one-earner household, but there is an effect on the intensive margin. There is a significant effect of anti-discrimination laws in the intra-household difference in hours worked for male same-sex and female same-sex households. An increase in the difference of hours worked within the household could suggest a greater specialization in market production for one person and a greater specialization in household production and working fewer hours for the other person in the partnership. Female same-sex households do see a significant increase in the likelihood of having any child and the number of children relative to male same-sex households.
This evidence suggests that men and women in same-sex partnerships are responding differently to these anti-discrimination laws. Female same-sex households begin to have more children than male same-sex households. It may appear in the main results that anti-discrimination laws do not help women in female same-sex partnerships. However, their response in the labor market could be driven by a greater desire for children and subsequent changes in household labor allocation due to an increase in job security. The corresponding changes in the labor market could represent an increase in welfare for female same-sex households if they value their new children more than their lost hours worked and subsequently lost wages. Consistent with the theory and empirical findings, female same-sex households could become more secure with one woman being the primary earner following anti-discrimination laws.

### v3

@@ 5.4 Metropolitan Areas
We re-estimate our model specifically for residents in metropolitan areas to assess whether the effects of anti-discrimination laws vary by geographic context. Table 8 shows that, within metro areas, the qualitative patterns for men in same-sex couples remain consistent with the main results: extensive-margin effects are positive and statistically significant, and the joint-estimation framework again shows that these effects are significantly larger for men than for women.
Table 8: Effects of Anti-Discrimination Laws on Labor Supply: Analysis of Metro Areas
(1) (2) (3) (4) (5) (6)
Extensive Margin Intensive Margin
Metro Areas
ENDA $\times$ SSP: Female -0.0262∗∗ -0.0284∗∗ -0.4601 -1.3804∗∗ -0.4479 -1.3355∗∗∗
(0.0115) (0.0129) (0.7244) (0.6511) (0.6693) (0.3471)
ENDA $\times$ SSP: Male 0.0352∗∗ 0.0398∗∗ 0.8139 1.4503 -0.4661 -0.2430
(0.0159) (0.0162) (0.5418) (0.7403) (0.3735) (0.3455)
$p$: M – F (Metro) 0.0005 0.0007 0.1692 0.0013 0.9819 0.0248
$p$: M (Metro – Non-Metro) 0.6882 0.5923 0.6851 0.2116 0.6688 0.0463
$p$: F (Metro – Non-Metro) 0.0263 0.0621 0.0065 0.1878 0.0019 0.1576
Observations 7,781,334 7,781,334 7,781,334 7,781,334 5,765,559 5,765,559
Mean (F) 0.834 0.801 35.20 40.42 41.43 47.52
Mean (M) 0.847 0.812 36.50 41.09 42.62 47.96
  • Notes: Data are derived from the American Community Survey (ACS) conducted annually between 2005 and 2019. We compare individuals in same-sex partnerships with those in different-sex partnerships within four years of ENDA enactment. All regressions in this table are estimated jointly with a fully interacted model that allows ENDA effects to vary across four groups: men and women, each in the metro subsample and in the non-metro subsample. This framework nests the separate-by-sex estimations and enables formal tests of equality across groups. We report the ENDA effects estimated for the metro subsample and include $p$-values testing whether these coefficients differ from those in the non-metro population. Columns (1–2) focus on the extensive margin, examining labor force participation and employment. Columns (3–6) analyze intensive-margin outcomes, usual weekly hours and weeks worked, with Columns (5–6) restricted to employed individuals. The “Mean (F)” and “Mean (M)” rows report pre-treatment averages of the outcome variables (including never-treated counties) for women and men in same-sex couples in the metro-area sample, respectively. Standard errors are clustered at the county level. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
For women in same-sex couples, the metro-area estimates become negative across several outcomes, including labor force participation, employment, and weeks worked. The joint model further indicates that the coefficients for women differ significantly between metro and non-metro samples, whereas for men no such differences are detected. One interpretation is that urban labor markets, combined with higher childcare costs, may amplify intra-household specialization, leading some women to reduce labor supply when partners experience greater employment stability. The metro-area analysis reinforces the gender differences observed in the main results: male extensive-margin gains remain robust, while female responses vary substantially by geography.
5.5 Residence - Work Mismatch
We next examine whether commuting patterns moderate the effects of anti-discrimination laws. In particular, we distinguish individuals who both live and work in the same county (“non-commuters”) from those who commute across county lines. Table 9 reports the estimates for non-commuters based on the fully interacted joint-estimation framework.
Table 9: Effects of Anti-Discrimination Laws on Labor Supply: Analysis of Work & Home Residence
(1) (2) (3) (4) (5) (6)
Extensive Margin Intensive Margin
Work–Residence County Matched (Same County)
ENDA $\times$ SSP: Female -0.0020 -0.0041 1.3169∗∗ -0.2497 1.1835∗∗∗ -0.6671
(0.0133) (0.0158) (0.5689) (0.7315) (0.3187) (0.4222)
ENDA $\times$ SSP: Male 0.0262∗∗ 0.0231 0.5135 1.0258 -0.4456 -0.1132
(0.0133) (0.0139) (0.5081) (0.6102) (0.3420) (0.3172)
$p$: M – F (Matched) 0.0858 0.2049 0.2584 0.1563 0.0011 0.2927
$p$: M (Matched – Mismatched) 0.0490 0.0960 0.0809 0.0271 0.2188 0.3734
$p$: F (Matched – Mismatched) 0.8780 0.7938 0.0155 0.3999 0.0291 0.5921
Observations 7,781,313 7,781,313 7,781,313 7,781,313 5,765,525 5,765,525
Mean (F) 0.779 0.736 32.70 37.46 40.87 46.77
Mean (M) 0.785 0.741 33.69 37.83 42.07 47.23
  • Notes: Data are derived from the American Community Survey (ACS) conducted annually between 2005 and 2019. We compare individuals in same-sex partnerships with those in different-sex partnerships within four years of ENDA enactment. All regressions in this table are estimated jointly with a fully interacted model that allows ENDA effects to vary across four groups: men and women, each in the commuter subsample and in the non-commuter subsample. This framework nests the separate-by-sex estimations and enables formal tests of equality across groups. We report ENDA effects for the matched sample (work county = residence county) and include p-values comparing these estimates to those obtained for individuals with a work–home county mismatch. Columns (1–2) focus on the extensive margin, examining labor force participation and employment. Columns (3–6) analyze intensive-margin outcomes, usual weekly hours and weeks worked, where Columns (5–6) restrict the sample to employed individuals. The “Mean (F)” and “Mean (M)” rows report pre-treatment averages of the outcome variables (including never-treated counties) for women and men in same-sex couples in the matched sample, respectively. Standard errors are clustered at the county level. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
Among non–commuters, the qualitative patterns for men in same–sex couples are broadly consistent with the main results: ENDA generates positive and statistically significant effects on labor force participation and employment (Columns 1–2), while intensive–margin effects are more modest. For women, extensive–margin coefficients remain close to zero, but we find sizable and statistically significant increases in usual weekly hours. The joint model shows that, within the non–commuter subsample, gender differences on the extensive margin are at most marginally significant, whereas women exhibit significantly larger ENDA effects than men on weekly hours among the employed. These marginal or insignificant gender differences likely reflect reduced precision due to the substantial sample restriction rather than substantive shifts in underlying patterns.
The pp–values comparing non–commuters to the commuter sample indicate that men’s ENDA effects change somewhat across commuting regimes on both the extensive and intensive margins, but the signs and overall pattern remain stable. For women, by contrast, the intensive–margin responses in hours differ more sharply between commuters and non–commuters.
@@ 5.6 Reporting Changes
Prior research has highlighted higher rates of misclassification errors before 2008, where different-sex couples were sometimes incorrectly identified as same-sex couples due to typos in the sex variable. In response, and considering the 2008 ACS redesign aimed at reducing such errors (o2009changes), we restrict our sample to 2008–2019 to test the robustness of our findings reported in Table 4. Table 10 presents the estimates based on this restricted sample, using the joint estimation framework.
Table 10: Effects of Anti-Discrimination Laws on Labor Supply: ACS Redesign (2008–2019)
(1) (2) (3) (4) (5) (6)
Extensive Margin Intensive Margin
ACS Redesign Sample (2008–2019)
ENDA $\times$ SSP: Female 0.0217 0.0272 1.3482∗∗ 0.6313 0.5888 -0.1525
(0.0129) (0.0195) (0.6583) (0.7446) (0.4430) (0.4733)
ENDA $\times$ SSP: Male 0.0459∗∗∗ 0.0492∗∗∗ 1.5759∗∗ 3.1075∗∗∗ -0.2548 0.8835∗∗∗
(0.0140) (0.0156) (0.7054) (0.8309) (0.7122) (0.3363)
$p$: M – F (ACS Redesign) 0.1635 0.3301 0.8272 0.0116 0.3877 0.0347
$p$: M (ACS Redesign – Rest) 0.7712 0.6221 0.5605 0.0837 0.8041 0.0165
$p$: F (ACS Redesign – Rest) 0.1277 0.0633 0.7380 0.1901 0.7515 0.0070
Observations 7,778,320 7,778,320 7,778,320 7,778,320 5,763,114 5,763,114
Mean (F) 0.824 0.787 34.27 39.72 41.09 47.51
Mean (M) 0.837 0.802 35.66 40.48 42.41 48.11
  • Notes: Data are derived from the American Community Survey (ACS) conducted between 2005 and 2019, following the major ACS redesign that altered survey implementation and imputation procedures. All regressions are estimated jointly using a fully interacted model that allows ENDA effects to vary across men and women within the redesign subsample. We report the ENDA effects estimated for the 2008–2019 sample and include $p$-values testing whether (i) male and female coefficients differ within the redesign sample, and (ii) redesign coefficients differ from those estimated in the 2005–2007 sample. Columns (1–2) report extensive-margin outcomes (labor force participation and employment). Columns (3–6) report intensive-margin outcomes (usual weekly hours and weeks worked), with Columns (5–6) restricted to employed individuals. The “Mean (F)” and “Mean (M)” rows report pre-treatment averages of the outcome variables (including never-treated counties) for women and men in same-sex couples in the ACS redesign sample (2008-2019), respectively. Standard errors are clustered at the county level. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
@@
For men in same-sex couples, the results remain qualitatively similar to our full-sample analysis: ENDA continues to generate positive and statistically significant effects on labor force participation and employment, with magnitudes close to those in Table 4. Among women, extensive-margin coefficients remain small, with the labor-force estimate marginally significant, and the employment estimate statistically insignificant. The intensive-margin results are broadly consistent with the full sample, though with some variation in magnitude.
The joint model also allows us to test whether the 2008–2019 coefficients differ from those estimated in the 2005–2007 sample. For men, most outcomes do not differ significantly across samples, though the effect on weeks worked does. For women, we detect significant differences in employment and in weeks worked among the employed. The male–female comparison on the extensive margin is not statistically significant in the ACS Redesign subsample, although the differences remain qualitatively similar to those in the full sample.
@@
Taken together, these results indicate that potential misclassification errors in earlier ACS waves do not drive our main conclusions. The consistency between the full and redesign-restricted samples reinforces the robustness of our empirical strategy and the validity of our conclusions regarding the impact of anti-discrimination laws on labor market outcomes for same-sex couples.
@@ 6.1 Differences in Response by Sex
The estimates reveal meaningful heterogeneity by sex within same-sex partnerships. For men in male same-sex couples, ENDA adoption significantly reduces the labor supply differential relative to men in different-sex partnerships. By contrast, we do not detect statistically significant average effects for women in female same-sex partnerships. Importantly, the absence of an average effect for women does not imply that ENDAs only affect men or that a single mechanism must operate symmetrically across groups. Our heterogeneity tests show that the extensive-margin responses for men and women differ significantly: men exhibit clear increases in labor supply, whereas the corresponding female estimates are near zero (and negative in metro areas). The appropriate interpretation is therefore one of heterogeneous adjustment, not opposing effects from a unified mechanism. Men respond on labor-market margins, while women appear to adjust through channels outside those margins, such as fertility decisions, with subsequent evidence also showing increased specialization for both male and female households.
A natural explanation draws on household models of specialization. If ENDAs reduce discrimination and increase job security for sexual minority workers, households may respond by reallocating labor between market and non-market activities. Prior work documents that reduced discrimination can strengthen the position of the primary earner within sexual minority households (badgett2006discrimination). In collective and bargaining models of the household, an improvement in one partner’s labor market opportunities increases the returns to specialization: the partner with stronger market prospects supplies more labor, while the other reallocates time toward household production (chiappori1992collective; lundberg1993separate).
Such mechanisms can generate heterogeneous responses across male and female same-sex households. Female same-sex couples are substantially more likely to have children than male same-sex couples, and thus may be more sensitive to changes in expected job stability or income that facilitate childcare. A shift toward specialization in female households may therefore involve adjustments in fertility and household production, rather than intensive-margin labor supply. To test these mechanisms directly, we estimate a household-level difference-in-differences model comparing male and female same-sex partnerships:
@@
$Y_{jct}$ $=\alpha\;+\;\beta_{1}\bigl[\mathrm{FemSSP}_{j}\times\mathrm{ENDA}_{ct}\bigr]\;+\;\beta_{2}\,\mathrm{FemSSP}_{j}\;+\;\beta_{3}\,\mathrm{ENDA}_{ct}+X_{jct}^{\prime}\delta+\mu_{c}+\delta_{t}+\epsilon_{jct}$ (3)
Here $\beta_{1}$ captures how female same-sex households respond relative to male same-sex households. We examine three sets of outcomes: (i) whether the household is a one-earner family, (ii) the absolute difference in weekly hours worked between partners (capturing intra-household specialization), and (iii) fertility outcomes.
Table 11 shows that ENDAs do not affect the probability of being a one-earner household, but they significantly increase intra-household differences in hours worked for both male and female same-sex households. The interaction term is negative because male households experience a larger increase in specialization; combining the ENDA main effect and the interaction yields a positive, though smaller, specialization response for female households, consistent with the specialization hypothesis.
Table 11: Responses between Male and Female Same-Sex Partnerships
(1) (2) (3) (4)
One Earner HH Diff in Hours Worked Any Children Number of Children
ENDA $\times$ Female 0.0107 -1.0573 0.0338∗∗∗ 0.0562
(0.0086) (0.5452) (0.0121) (0.0303)
ENDA -0.0201 1.6692∗∗∗ -0.0056 -0.0289
(0.0133) (0.4773) (0.0148) (0.0329)
Female 0.0029 -0.8066∗∗∗ 0.1428∗∗∗ 0.2273∗∗∗
(0.0061) (0.2587) (0.0068) (0.0157)
Observations 41,382 41,382 41,382 41,382
Mean (F) 0.2585 26.8599 0.3460 0.6036
$R^{2}$ 0.0534 0.1349 0.1306 0.1216
  • Notes: Data is derived from the Annual Community Surveys (ACS) spanning 2005-2019 comparing women in same-sex partnerships to men. The first column examines if the household has only one earner. The second column tests the difference in absolute terms of hours worked between the two partners. The third and fourth columns examine how households differ in having children. The coefficients show the differential effect of anti-discrimination between female and male same-sex households. The “Mean (F)” row reports pre-treatment averages of the outcome variables (including never-treated counties) for female same-sex couple households. Standard errors are clustered at the county level. $p<0.10$, ∗∗ $p<0.05$, $^{{}^{***}}$ $p<0.01$.
Female same-sex households also exhibit significant increases in fertility, both in the likelihood of having any child and in the number of children, relative to male households. These results suggest that ENDA-induced improvements in job security may allow female same-sex households to realize fertility preferences and subsequently adjust labor supply to accommodate additional childcare needs.
Taken together, these findings show that male and female same-sex households respond differently to anti-discrimination laws not because the laws are ineffective for women, but because the relevant channels of adjustment differ. Male same-sex households respond primarily through increased labor-market specialization, while female same-sex households adjust fertility and household production. If female households value these additional children more than forgone labor market hours, these responses may reflect higher welfare even in the absence of average improvements in women’s labor market outcomes.
6.2 Sorting and Partnership Formation
A remaining concern is whether ENDA adoption changes who appears in the same-sex partnered sample. If ENDAs affect partnership formation, cohabitation decisions, or relationship stability, then the composition of same-sex partnerships could shift endogenously with treatment. Such changes would complicate interpretation of the triple-differences estimates, since the observed sample would no longer represent the same underlying population before and after ENDA adoption. To assess this possibility, we examine whether the probability of being observed in a same-sex partnership changes following ENDA adoption, and we additionally test for sorting on observable demographics. Specifically we estimate:
$\mathrm{Y}_{ict}$ $=\beta\,\mathrm{ENDA}_{ct}+\mu_{c}+\delta_{t}+\epsilon_{ict},$ (4)
where $\mathrm{Y}_{ict}$ is an indicator for living in a same-sex partnership, age, or years of education. $\mu_{c}$ are county fixed effects, and $\delta_{t}$ are year fixed effects. Our core specification tests whether the fraction of adults observed in same-sex partnerships shifts systematically in treated counties relative to untreated counties. Using age and years of education as outcomes provides a direct check for demographic sorting: if ENDAs systematically attract (or repel) younger or more educated individuals, we would observe corresponding shifts in treated counties.
Table 12 reports the results. Across all specifications, and separately for men and women, we find no evidence that ENDAs affect the likelihood of living in a same-sex partnership. Coefficients are small, precisely estimated, and statistically insignificant. These findings indicate that ENDA adoption does not meaningfully alter partnership formation or cohabitation patterns.
Table 12: Sorting and Reporting
(1) (2) (3) (4) (5)
All SSP Male SSP Female SSP Age Years of Education
ENDA -0.0002 0.0004 -0.0008 0.0257 -0.0212
(0.0005) (0.0005) (0.0005) (0.0648) (0.0182)
Mean 0.0093 0.0087 0.0099 45.5421 13.6441
$R^{2}$ 0.0043 0.0081 0.0024 0.0101 0.0467
  • Notes: Data comes from the 2005-2019 yearly ACS seeing how the number of same-sex partnerships change in a county after the passage of an anti-discrimination law. The first column looks at both men and women with the next two columns separating the sexes. Column (4) investigates age and (5) looks at years of education. The “Mean” row reports pre-treatment averages of the outcome variables (including never-treated counties) for same-sex couples. Standard errors are clustered at the county level. $p<0.10$, ∗∗ $p<0.05$, $^{{}^{***}}$ $p<0.01$.
The absence of sorting responses strengthens the interpretation of the main results. Because ENDAs do not appear to change who enters or exits same-sex partnerships, the coupled-sample restriction does not generate treatment-induced compositional bias. Thus, the heterogeneous labor supply responses documented above are best understood as behavioral adjustments within stable households, rather than as artifacts of changes in relationship structure or partner selection.